5. A firm is issuing new debt to finance a capital investment project. The firm will issue 15,550 new bonds with a $1,000 face value that will mature in 10 years. The bonds will pay a $35 semiannual coupon, and similar bonds are currently priced at 95% of par. The associated flotation costs are expected to be $15 per bond. Further, the company has a marginal tax rate of 34%. Given this information, what is the before-tax cost of debt?

A)6.85%

B)3.98%

C)9.54%

D)8.15%

E)7.96%
Answer #1

Answer = E) 7.96

Explanation:

FV = -1000 PMT = -35 PV = 950-15 = 935, N = 10×2 = 20

I/Y = RATE(nper,pmt,pv,fv)

compute I/Y = 3.9772 × 2 = 7.955%